OpenStudy (haleyelizabeth2017):

Simplify the expression. Cos x+sin x tan x. Please no direct answers :)

OpenStudy (waleed_imtiaz):

you can simplify tanx and it can be written as Sinx/cosx; then Take the LCM

OpenStudy (waleed_imtiaz):

\[[\sin ^{2}x+\cos ^{2}x]/ \cos x\]